OpenStudy (anonymous):

I need to find the zeroes of f(x) = 5x2 + 2x + 1

OpenStudy (anonymous):

@AliceCullen @ramseysa @Elsa213 @DaBest21 @thomaster

OpenStudy (bloomlocke367):

zeroes are just your solutions. It's where x=0, or where the graph crosses the x-axis. Do you know how to solve quadratics?

OpenStudy (anonymous):

No @BloomLocke367

OpenStudy (bloomlocke367):

Here, take a look at this tutorial to see if you understand. If you still need help, please do tag me and I'll try my best to help you! http://openstudy.com/users/bloomlocke367#/updates/55acffcee4b071e6530c96ce

OpenStudy (anonymous):

Wait, is -1/5 and 4/5

OpenStudy (anonymous):

Are those the zeroes?

OpenStudy (bloomlocke367):

Actually, I'm pretty sure the zeroes are complex roots. One moment. I have to double check my work.

OpenStudy (anonymous):

- 1/5 and - 4/5

OpenStudy (bloomlocke367):

No, those are not the roots. They are, in fact, complex. Do you know what a complex number is?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

- 1/5 and - 2i/5

OpenStudy (bloomlocke367):

Basically, a complex number is a number that has an imaginary number in it. The often have a coefficient, i, and a constant. Here's an example of what it would look like: 3i+8. That's just an example, not the answer. Do you understand imaginary numbers? Where did you get those answers?

OpenStudy (anonymous):

Mathway, I suck at math and despise it. I'm sorry, I'm a history and English guy.

OpenStudy (anonymous):

Actually, I got them from Wolfram/Alpha

OpenStudy (bloomlocke367):

Really? huh. Try using the quadratic formula, which is \(\LARGE x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (bloomlocke367):

In this case, a is 5, b is 2, and c is 1.

OpenStudy (anonymous):

(-2±√(2^2-4*5*1))/(2*5)

OpenStudy (anonymous):

Is it 1/10(-2 ± 4i)?

OpenStudy (anonymous):

@badmood

OpenStudy (radar):

Actually you were right. You stated you had to find the 0"s of f(x) = 5x^2 + 2x + 1 This is the value where f(x) = 0. Obviously f(x) depends on x and x can be an infinite number of values both positive and negative and complex. When x = 0, in this given function f(x) = 1, when x = -1, f(x) = 4 etc. for f(x) = 0 you would solve for x, and the quadratic equation would be a reasonable method, however when you graph this function you will find that the f(x) nevers gets to 0. This does not mean that the function does not exists it simply means there is nor real "0" for f(x) for any real value of x.